OpenStudy (ashsmashum1290):

also accsidently pressed grad XD WILL FAN AND MEDAL PLZZZZ HELPPP !!!!!!which equation is true for all the values in the table? X:0,10,20,30 Y:-3,17,37,57

OpenStudy (ashsmashum1290):

A.Y=1/2x - 3 B.Y=5/7x - 3 C.Y=2x - 3 D.Y-x - 2

OpenStudy (ashsmashum1290):

@kropot72 @ShadowLegendX @whpalmer4

OpenStudy (ashsmashum1290):

@RedNeckOutLaw @Redcan @Miracrown @poopsiedoodle @Luigi0210

OpenStudy (ashsmashum1290):

@Miracrown

OpenStudy (anonymous):

try hint and trial method put values of x and find if it satisfies value of y in which equation

OpenStudy (whpalmer4):

or you could just work out the equation from two data points. if you know two points \((x_1,y_1), (x_2,y_2)\) then the slope of the line between them is given by \[m =\frac{y_2-y_1}{x_2-x_1}\] knowing the slope and a point you can get the equation of the line from \[y-y_1 = m(x-x_1)\] solve the equation for \(y\) to make it look like your answer choices took longer to type the explanation than to find the answer...

OpenStudy (ashsmashum1290):

um A??

OpenStudy (victoriasushchik):

it would be c

OpenStudy (whpalmer4):

no, A is not correct. if you show me your work, maybe we can figure out where you went wrong

OpenStudy (ashsmashum1290):

srry if im wrong im horrable at math

OpenStudy (whpalmer4):

no apology necessary, let's see if there's something we can help you do better.

OpenStudy (whpalmer4):

what did you find for the value of the slope (\(m\))?

OpenStudy (ashsmashum1290):

nvm gtg but thx :( i wanted to check this answer before i went to bed but dont have time :/ byw thx

OpenStudy (ashsmashum1290):

bye*

OpenStudy (whpalmer4):

ok